### v2

@@ Abstract
This study aimed to investigate the application of label propagation techniques to propagate labels among photoplethysmogram (PPG) signals, particularly in imbalanced class scenarios and limited data availability scenarios, where clean PPG samples are significantly outnumbered by artifact-contaminated samples. Our research compares the performance of supervised classifiers, such as conventional classifiers and neural networks (Multi-Layer Perceptron (MLP), Transformers, Fully Convolutional Network (FCN)), with the semi-supervised label propagation algorithm for artifact classification in PPG signals. The results indicate that the label propagation algorithm achieves a precision of 91%, a recall of 90%, and an F1 score of 90% for the ”artifacts” class, showcasing its effectiveness in annotating a medical dataset, even in cases where clean samples are rare. Although the K-Nearest Neighbors (KNN) supervised model demonstrated good results with a precision of 89%, a recall of 95%, and an F1 score of 92%, the semi-supervised algorithm excels in artifact detection. In the case of imbalanced and limited pediatric intensive care environment data, the semi-supervised label propagation algorithm is promising for artifact detection in PPG signals. The results of this study are important for improving the accuracy of PPG-based health monitoring, particularly in situations in which motion artifacts pose challenges to data interpretation.
@@ I Introduction
Machine learning, a sub-field of artificial intelligence [1], has emerged as a transformative technology in various domains, including healthcare. With its ability to analyze large amounts of data [2] and find important patterns, machine learning is a robust technique. It has the potential to improve healthcare outcomes, help doctors make better decisions [3], and revolutionize medical research. Using advanced algorithms and computing power, machine learning can extract all the necessary information from various types of healthcare data, such as electronic medical records [4], medical images, and physiological signals.
In clinical context, machine learning techniques can improve decision making. Support systems help with things like medication alerts and reminders for patient safety. Indeed, the model uses patient medical records and the vast amount of information available to develop models that aim to predict injuries [5], detect heart disease earlier [6] and mortality [7]. Based on clinical experience and physiological principles, machine learning provides scores that enable hospital staff to make faster, more accurate decisions. Additionally, machine learning algorithms can contribute to drug discovery and development by expediting the identification of promising drug targets, optimizing drug efficacy, and predicting potential adverse reactions [8].
Despite its potential, the integration of machine learning into healthcare comes with challenges and considerations. Privacy and ethical implications must be taken into account when applying machine learning to healthcare [9]. The data acquired must respect patient privacy and confidentiality. This data is often recorded by several devices, so it is important to link the recordings so that the data is standardized and can be more easily managed. In order to be able to link these data, they need to be collected in a single database, and thus be consistent with each other. In other words, the data must be harmonized [10]. Data security and accountability should be prioritized, especially when using cloud-based technologies that require extensive storage and computing power.
One major concern is the availability of high-quality data for training and testing these algorithms [11]. To be able to evaluate the performance of the algorithms implemented, it is necessary to have access to a ground truth. This ground truth is often non-existent or requires the help of experts, and is often time-consuming. Moreover, it’s crucial to have a large enough sample size to ensure accurate and reliable predictions. However, obtaining such labeled data can be challenging, as it may be limited and incomplete. Furthermore, the imbalance in classes during the algorithm’s training phase is a common consequence of medical data predominantly representing the most frequent cases present in patients. Consequently, a significant challenge arises in rebalancing the data whereas ensuring the retention of its crucial medical value. Additionally, the amount of erroneous, missing or imprecise data is an obstacle to model prediction. Erroneous data are data which have been altered by abnormal occurrences. Continuous waveform data captured by sensors is vulnerable to high-frequency artifacts that often result from patient motion or clinical interventions. Even periodic clinical measurements are not immune to contamination, as data collection and coding practices can introduce issues [10].
@@
The main objective of this work is to detect motion artifacts in photoplethysmogram (PPG) signals obtained from the pediatric intensive care unit (PICU) database of the CHU Sainte-Justine hospital (CHUSJ). PPG signals are frequently captured during different types of movements which introduces motion noise and interferes with the accuracy of the signals. This noise is irregular and causes high-amplitude fluctuations within the PPG signals [12]. Motion artifacts can result in the pulse oximeter either misinterpreting movement as the actual signal or masking the true signal with unwanted interference, which can lead to incorrect readings, false alarms, and missed important alarms [13]. The cleaned PPG signals will be used in the construction of clinical decision systems (CDSS) at the PICU of CHUSJ. Specifically, annotated signals will be used in screening and identifying various health-related concerns on children. As an example, changes in blood pressure in children serve as significant indicators for identifying patients who require immediate care and admission to the PICU. Invasive methods, like catheter insertion for continuous blood pressure monitoring, offer precise real-time data but come with significant risks such as bleeding and infection [14]. On the other hand, conventional cuff-based measurements, though less invasive, provide only intermittent readings and may not capture sudden clinical changes effectively. Therefore, predicting blood pressure from PPG waveforms has emerged as a successful approach [15] for comprehensive CDSS applications. In this project, we used a semi-supervised labeling algorithm to annotate PPG signals to effectively remove artifacts from them in real-time. A small proportion of data was annotated by an expert and a statistical analysis algorithm was then used to validate the annotations. To annotate the entire dataset, a machine learning algorithm that annotates all data using only a small proportion of previously annotated data, was used. The Label Propagation (LP) algorithm is a machine learning technique used for semi-supervised learning tasks. This algorithm is effective in our scenario because the availability of labeled data is limited and there is a large amount of unlabeled data [16].
Our medical data are very unbalanced, with around 80% of pulses free of artifacts and only 20% with artifacts. This means that to have an accurate labeling algorithm, a rebalancing of the classes in the training part need to be done. Several methods are available for this: oversampling, undersampling and, both oversampling and undersampling. It must be remembered that medical data is being worked with, so using sampling methods must make medical sense, whether by randomly duplicating data or by removing it. Medical data involves intricate relationships among different data elements, such as patient demographics, medical history, symptoms, diagnoses, treatments, and outcomes [17].
@@
Another aim of this project is to compare classifiers to the Label Propagation algorithm, used as a classifier, to accurately detect artifacts. In [18] the authors explore the use of semi-supervised models to classify temporal data. These models are based on a graphical approach, such as the label propagation algorithm. The results of the algorithm are evaluated on different datasets of varying lengths, including ECG (electrocardiogram) signal data. The results show that semi-supervised models are accurate for classifying time series data. However, these algorithms have not been applied to artifact detection. Semi-supervised learning is widely used as a classification algorithm in cases where not all data is annotated. Active learning is also a very powerful semi-supervised classification method that has proven its effectiveness for temporal data [19].
@@
Machine learning classifiers are designed to categorize data into distinct classes or groups based on specific features or patterns. The use of classifiers has witnessed significant growth across different domains, especially in health care, where classifiers are a real help in decision-making [20]. By leveraging statistical analysis and optimization techniques, classifiers enable accurate predictions and information retrieval. The spectrum of classifiers is very wide: from traditional classifiers like K-Nearest Neighbors (KNN), Support Vector Machine (SVM), Decision Tree (DT), Naive Bayes classifier (NB) [21], to classifiers using neural networks, such as Multi-Layer Perceptron classifier (MLP) or Transformers. A comparison of the results of each type of classifier with the semi-supervised Label Propagation algorithm will be dressed. The effectiveness of these two streams is analyzed by the experimental results (in section IV) from the comparative analysis of semi-supervised label propagation (label propagation with KNN kernel) and fully-supervised learning, including conventional machine learning classifiers (KNN, Support Vector Classification (SVC), DT, Random Forest (RF), GaussianNB, MultinominalNB, and Logistic Regression (LR)), MLP, and Transformers. Then, the best classification method will be presented, followed by a conclusion on artifact detection.
@@
The paper is structured as follows. In section II, data characteristics, preprocessing and methodology: labeling and classification, are introduced. In section III, the implementation of experiments is presented. Section IV is used to evaluate the results with different metrics and present a comparison of experimental result tables. In section V, the results are interpreted and the limitations are discussed.
II Material and Methods
This study was conducted following ethical approval from the research ethics board at CHUSJ (protocol number 2023-4556, accepted January 18, 2023). The detailed workflow of the various work stages is shown in the Fig. 1.
II-A Data Collection
@@
The aim of this project is to detect motion artifacts in PPG signals. The eligible study population includes all children aged 0 to 18 years, admitted between September 2018 and July 2022 inclusive, for whom electrocardiogram (ECG), PPG and arterial blood pressure (ABP) waveform records are available. In this population, specific exclusion criteria have been established to avoid bias. Data collected beyond the fourth day of hospital stay will be disregarded to prevent potential bias from a few patients who may have prolonged stays with arterial lines. Patients on extracorporeal membrane oxygenation (ECMO) treatment will also be excluded from the analysis. Furthermore, if a patient is readmitted to the PICU multiple times, only data from the first stay will be considered for analysis.
@@
A PPG signal is recorded using a sensor called the pulse oximeter. This device is placed on a patient’s skin, for children on a fingertip or earlobe. A PPG sensor emits light into the skin, which is partially absorbed by the blood vessels. Changes in blood flow during the cardiac cycle cause variations in light absorption. The sensor detects the reflected light, measuring its intensity modulated by blood volume changes. This varying intensity is converted into an electrical signal, creating the PPG waveform. Blood pressure signals are recorded using an invasive and continuous method, i.e., the catheter, and a non-invasive and discontinuous method, i.e., the blood pressure cuff. ECG is continuously recorded by placing electrodes on the patient’s chest. The Sainte-Justine University Hospital Pediatric Intensive Care Unit (PICU) utilizes a high-resolution research database (HRDB) [22, 23] that has been approved by the ethical committee. The HRDB links biomedical signals extracted from the different devices, displayed through patient monitors, to the electronic patient record continuously throughout their stay in the unit [24].
Between 2018 and 2022, 1571 patients met the inclusion criteria. For each patient, four physiological signals were extracted: ECG, PPG, blood pressure from the catheter, and blood pressure from the cuff. Each signal was extracted over a period of 96 hours (4 days). Signal values are grouped together in a table with the date and time of acquisition. For the PPG signal, 640 values are acquired every 5 seconds, corresponding to a sampling frequency of 128 Hz. For blood pressure and ECG signals, 2560 values are acquired every 5 seconds, with a sampling frequency of 512 Hz. For the duration of the extraction, a fixed 30-second window of PPG signals will be used for the further processing.
II-B Preprocessing
@@
    Filtering: each signal window is filtered using a band-pass Butterworth filter, the cut-off frequencies are 0.5 and 5 Hz, corresponding to a heart rate between 30 and 300 bpm. A forward-backward filtering is used to avoid phase distortions. The objective is to remove baseline wander and high-frequency noise.
@@
    Pulse segmentation: a function to finds all local minima by a comparison between samples is used. The aim is to divide the preprocessed PPG signal into smaller segments or windows to be able to detect the artifacts present for each pulse. In our case, a segment is a pulse. The size of each segment may vary depending on the characteristics of the PPG signal and the specific application the signal pulses. A pulse is considered to lie between two minima.
@@
    Resampling: the duration of a cardiac cycle for children is between 0.3 and 1 second. A pulse is representing a cardiac cycle. Therefore, not all pulses have the same number of samples. Each pulse is uniformly oversampled in time to contain 256 samples, which corresponds to a heart cycle of 1s. A linear interpolation function [26] is used to create the missing points for each pulse. Linear interpolation is favored for signals due to its simplicity, computational efficiency, and ability to estimate values between known data points. It maintains signal continuity and linearity, making it suitable for signals with relatively smooth and linear variations.
@@
    Data transformation: each PPG pulse, which is essentially a waveform representing blood volume changes over time, can be represented as a data point in a column that contains 256 values. These values are equally spaced points obtained using step 3 of the preprocessing. At the end of preprocessing, a vector of 256 points is obtained, representing a pulse of the PPG signal. The number of vectors depends on the number of pulses. This method allows to work with PPG data in a structured manner, suitable for a range of applications, from statistical analysis to machine learning.
Figure 2: Example of a 10s segment of a 30s raw PPG signal in the top image, filtered signal in the middle image and segmented signal in the bottom image.
The Fig. 2 shows the first 10 seconds of a raw PPG signal, then when the signal has been filtered and finally when the pulses have been segmented. The effect of the bandpass filter can be seen in the second figure. The filter has smoothed the signal by removing the extreme frequency components. The signal waveform is preserved, and the filter does not introduce resonances or significant ripples in the desired frequency range. Note that the first pulse has not been segmented. This is because the function could not detect the two minima that make up a pulse and therefore could not segment it. The signal does not start at the first low point of the pulse.
@@
II-C Dataset annotation
First, after preprocessing the data, the aim is to build a ground truth for future evaluation of the classification algorithms. In this objective, one human expert annotated 10% of the database. Annotation is visual, comparing pulses with each other and binary classifying each pulse as good or as containing artifacts. To avoid involving another human expert, in consideration of time and specialist resources, we implemented an automated algorithm to handle additional annotations. This algorithm, acting as a surrogate expert, was developed to recheck the entirety of the 10% annotated data by the human expert, identifying similarities in the process. Employing a statistical approach, the algorithm determines if the values of a given pulse lie within standard parameters or deviate from the norm. We subsequently cross-validated the algorithm’s annotations against those from the professional expert to determine the algorithm’s accuracy. It was decided to annotate a maximum of 10% of the database and then use the label propagation algorithm to annotate the rest of the data.
II-C1 Expert labeling
PPG signals already segmented are presented to the expert. By analyzing each pulse, the expert classifies each pulse as artifact or artifact-free. A pulse is defined as artifact-free if its morphology is typical i.e. if its characteristics - amplitude, width, shape - are the same as those of adjacent signals. A pulse is defined with artifacts if its characteristics differ from those of adjacent pulses (see Fig. 3). To recheck the annotations of one expert, an algorithm that acts as a second expert was set up, allowing all impulses to be reannotated to see similarities. This algorithm uses a statistical approach to assess whether the statistical values of a pulse are normal or outside the norm.
II-C2 Statistical analysis
For each cardiac cycle, which corresponds to a pulse, if the waveform is similar, statistics such as skewness, standard deviation (std) and kurtosis are approximately constant for each cycle. It is therefore possible to detect motion artifacts by using the value of these statistics to differentiate a pulse without artifacts from a pulse with motion artifacts [27]. Skewness indicates the degree of asymmetry in the probability distribution of a random variable around its mean. It can take on positive, zero, negative, or undefined values, reflecting the shape and symmetry of the distribution. Kurtosis is the sharpennes of the peak of a frequency-distribution curve and standard deviation reflects the dispersion degree of a data set. If $X$ is considered a variable with $\mu$ and $\sigma$ the mean and standard deviation respectively, statistical values are calculated as follows:

### v3

@@ Abstract
This study aimed to investigate the application of label propagation techniques to propagate labels among photoplethysmogram (PPG) signals, particularly in imbalanced class scenarios and limited data availability scenarios, where clean PPG samples are significantly outnumbered by artifact-contaminated samples. We investigated a dataset comprising PPG recordings from 1571 patients, wherein approximately 82% of the samples were identified as clean, while the remaining 18% were contaminated by artifacts. Our research compares the performance of supervised classifiers, such as conventional classifiers and neural networks (Multi-Layer Perceptron (MLP), Transformers, Fully Convolutional Network (FCN)), with the semi-supervised Label Propagation (LP) algorithm for artifact classification in PPG signals. The results indicate that the LP algorithm achieves a precision of 91%, a recall of 90%, and an F1 score of 90% for the ”artifacts” class, showcasing its effectiveness in annotating a medical dataset, even in cases where clean samples are rare. Although the K-Nearest Neighbors (KNN) supervised model demonstrated good results with a precision of 89%, a recall of 95%, and an F1 score of 92%, the semi-supervised algorithm excels in artifact detection. In the case of imbalanced and limited pediatric intensive care environment data, the semi-supervised LP algorithm is promising for artifact detection in PPG signals. The results of this study are important for improving the accuracy of PPG-based health monitoring, particularly in situations in which motion artifacts pose challenges to data interpretation.
@@ I Introduction
Machine learning, a sub-field of artificial intelligence [1], has emerged as a transformative technology in various domains, including healthcare. With its ability to analyze large amounts of data [2], it has the potential to improve healthcare outcomes, help doctors make better decisions [3], and revolutionize medical research with models that aim to predict injuries [4], detect heart disease earlier [5] and mortality [6]. Additionally, machine learning algorithms can contribute to drug discovery and development, optimizing drug efficacy and predicting potential adverse reactions [7]. Machine learning can extract all the necessary information from various types of healthcare data, such as electronic medical records [8], medical images, and physiological signals.
Despite its potential, the integration of machine learning into healthcare comes with challenges and considerations. Privacy and ethical implications must be taken into account [9]. The data acquired must respect patient privacy and confidentiality and also require standardization and centralized collection for ease of management and consistency, ensuring harmonization [10]. One major concern is the availability of high-quality data for training and testing these algorithms [11]. To evaluate the performance of the algorithms implemented, it is necessary to have access to a ground truth. Accessing ground truth for evaluating algorithms is challenging, often requiring expert input and large, complete datasets, particularly due to class imbalances in medical data. This further complicates model training, necessitating rebalancing while preserving medical value, with erroneous, missing, or imprecise data exacerbated by artifacts from patient motion or clinical interventions posing additional obstacles to accurate predictions.
@@
During a patient’s stay in the hospital, it is important to constantly monitor vital signs. One of these vital signals is the PPG signal, which is frequently captured during different types of movements, introducing motion noise and interfering with the accuracy of the signals. This noise is irregular and causes high-amplitude fluctuations within the PPG signals [12]. Motion artifacts can result in the pulse oximeter either misinterpreting movement as the actual signal or masking the true signal with unwanted interference, leading to incorrect readings, false alarms, and missed important alarms [13]. The main objective of this work is to detect motion artifacts in PPG signals obtained from the Pediatric Intensive Care Unit (PICU) database of the CHU Sainte-Justine Hospital (CHUSJ). The cleaned PPG signals will be used to construct clinical decision systems (CDSS) at CHUSJ’s PICU. Specifically, annotated signals will be used in screening and identifying various health-related concerns in children. For example, changes in blood pressure in children are significant indicators for identifying patients who require immediate care and admission to the PICU. Invasive methods, like catheter insertion for continuous blood pressure monitoring, offer precise real-time data but come with significant risks such as bleeding and infection [14]. On the other hand, conventional cuff-based measurements, though less invasive, provide only intermittent readings and may not capture sudden clinical changes effectively. Therefore, predicting blood pressure from PPG waveforms has emerged as a successful approach [15] for comprehensive CDSS applications.
This study contributes to the field in three main ways. Firstly, we compare resampling methods commonly used in medical data analysis to address the imbalance between clean PPG samples and artifact-contaminated ones. Secondly, we validate the efficacy of the LP algorithm for motion artifact detection within PPG signals, offering insights into its performance in scenarios with limited labeled data. Lastly, we present a detailed performance comparison between traditional supervised algorithms and the semi-supervised LP approach, highlighting the advantages of leveraging unlabeled data in artifact classification tasks.
II Related work
Numerous methods have already been developed to detect motion artifacts in PPG signals. First, the traditional methods are easy to implement. In [16], the authors used statistical analysis to compare the values of three statistics calculated for each pulse of the PPG signal to determine which pulses are noisy. This method will be used in the labeling step for the rest of the project. Adaptive filtering is another method of artifact detection [17]. The adaptive filter uses an algorithm that continuously updates its coefficients to obtain an error signal as close as possible to the original PPG signal. Both approaches have the advantage of being easy to implement but are notably sensitive to empirical thresholds. Among other popular methods, the wavelet transform uses cascaded high-pass and low-pass filters to obtain the desired signal decomposition. Once the signal has been decomposed, the coefficients are analyzed to identify any artifacts [18]. Empirical mode decomposition, like the wavelet transform, is a time-frequency analysis of the signal [19]. When these modes are obtained, the objective is to calculate the instantaneous frequency for each mode to detect modes that have a frequency close to the harmonics of a PPG signal and modes characteristic of motion artifacts. So, these methods have the advantage of being fast and simple, which is useful, but when used alone, they have limited adaptability and may not work as well with complex movements or unexpected scenarios. A summary table with the review activity is presented in Table I. Therefore, we decided to use a combination of signal processing algorithms for the preprocessing part and machine learning models.
@@ II Related work
Regarding machine learning models, in [20], the authors explore using semi-supervised models to classify temporal data. These models are based on a graphical approach like the LP algorithm. The algorithm’s results are evaluated on different datasets of varying lengths, including ECG (electrocardiogram) signal data. The results show that semi-supervised models are accurate for classifying time series data. However, these algorithms have not been applied to artifact detection. Semi-supervised learning is widely used as a classification algorithm in cases where not all data is annotated. Active learning is also a powerful semi-supervised classification method that has proven effective for temporal data [21]. In our scenario, the LP algorithm is effective because the availability of labeled data is limited, and there is a large amount of unlabeled data [22]. Considering this information, the LP algorithm was implemented for this project.
In the first part of the project, the LP algorithm is used for data annotation. First, an expert annotated a small proportion of data, and a statistical analysis algorithm was used to validate the annotations. Then, the LP annotates all data using only a small proportion of previously annotated data. Our medical data are unbalanced, with around 80% of pulses free of artifacts and only 20% with artifacts. This means that to have an accurate labeling algorithm, a rebalancing of the classes in the training part needs to be done. Several methods are available for this: oversampling, undersampling, and both oversampling and undersampling. It must be remembered that medical data is being worked with, so sampling methods must make medical sense, whether by randomly duplicating data or by removing it. Medical data involves intricate relationships among data elements, such as patient demographics, medical history, symptoms, diagnoses, treatments, and outcomes [23].
TABLE I: Summary table of the literature review
Authors Title Year Source Findings
Q. Wang et al. Artifact reduction based on empirical mode decomposition (EMD) in photoplethysmography for pulse rate detection 2010 2010 Annual International Conference of the IEEE Engineering in Medicine and Biology Empirical Mode Decomposition and Hilbert transform combined give good results for the decomposition of PPG signals and the reduction of motion artifacts
G. Joseph et al. Photoplethysmogram (PPG) signal analysis and wavelet de-noising 2014 2014 Annual International Conference on Emerging Research Areas: Magnetics, Machines and Drives (AICERA/iCMMD) Unwanted PPG signal interference is successfully removed using wavelet transform while preserving the signal information
Z. Xu and K. Funaya Time series analysis with graph-based semi-supervised learning 2015 2015 IEEE International Conference on Data Science and Advanced Analytics (DSAA) The use of a new probabilistic semi-supervised method combining different graph constructions and distance techniques gives better results on different types of real data
S. Hanyu and C. Xiaohui Motion artifact detection and reduction in PPG signals based on statistics analysis 2017 29th Chinese Control And Decision Conference (CCDC) The use of statistical thresholds on corrupted PPG segments correlated with high-quality segments effectively removes motion artifacts
C.-C. Wu et al. An implementation of motion artifacts elimination for PPG signal processing based on recursive least squares adaptive filter 2017 2017 IEEE Biomedical Circuits and Systems Conference (BioCAS) Adaptive approach to remove motion artifacts, using DC Remover method and Recursive Least Squares adaptive filter
Y. Shin et al. Coherence-based label propagation over time series for accelerated active learning 2021 International Conference on Learning Representations A new active learning method for time series, called TCLP, improves classification accuracy when a very small number of data points are already labeled
D. Bünger et al. An empirical study of graph-based approaches for semi-supervised time series classification 2022 Frontiers in Applied Mathematics and Statistics, vol. 7 Comparison of different distance measures in the implementation of graph-based models, including some semi-supervised models, in classifying binary time series datasets
@@
Another aim of this project is to compare classifiers to the LP algorithm, used as a classifier, to accurately detect artifacts. In health care, classifiers are a real help in decision-making [24]. The spectrum of classifiers is very wide: from traditional classifiers like KNN, Support Vector Machine (SVM), Decision Tree (DT), and Naive Bayes classifier (NB) [25], to classifiers using neural networks, such as MLP or Transformers. A comparison of the results of each type of classifier with the semi-supervised LP algorithm will be presented. The effectiveness of these two streams is analyzed by the experimental results (in section V) from the comparative analysis of semi-supervised LP (with KNN kernel) and fully-supervised learning, including conventional machine learning classifiers (KNN, Support Vector Classification (SVC), DT, Random Forest (RF), GaussianNB, MultinominalNB, and Logistic Regression (LR)), MLP, and Transformers. Then, the best classification method will be presented, followed by a conclusion on artifact detection.
@@
The paper is structured as follows. In section III, data characteristics, preprocessing, methodology, labeling, and classification are introduced. In section IV, the implementation of experiments is presented. Section V is used to evaluate the results with different metrics and present a comparison of experimental result tables. In section VII, the results are interpreted, and the limitations are discussed.
III Material and Methods
This study was conducted following ethical approval from the research ethics board at CHUSJ (protocol number 2023-4556, accepted January 18, 2023). The detailed workflow of the various work stages is shown in Fig. 1. Specifically, the workflow of the proposed method for detecting motion artifacts in PPG signals begins with the input of a 30-second PPG signal. This signal undergoes data preprocessing, which includes filtering, segmentation, resampling, and normalization to prepare the data for analysis. Next, a labeling and classification step is performed using a label propagation algorithm to identify and classify segments of the signal. Finally, the process outputs artifact detection, highlighting the portions of the signal affected by motion artifacts.
III-A Data Collection
@@
This project aims to detect motion artifacts in PPG signals. The eligible study population includes all children aged 0 to 18 years, admitted between September 2018 and July 2022 inclusive, for whom electrocardiogram (ECG), PPG, and arterial blood pressure (ABP) waveform records are available. In this population, specific exclusion criteria have been established to avoid bias. Data collected beyond the fourth day of hospital stay will be disregarded to prevent potential bias from a few patients who may have prolonged stays with arterial lines. Patients on extracorporeal membrane oxygenation (ECMO) treatment will also be excluded from the analysis. Furthermore, if a patient is readmitted to the PICU multiple times, only data from the first stay will be analyzed.
@@
A PPG signal is recorded using a sensor called the pulse oximeter. This device is placed on a patient’s skin, for children, on a fingertip or earlobe. A PPG sensor emits light into the skin, partially absorbed by the blood vessels. Changes in blood flow during the cardiac cycle cause variations in light absorption. The sensor detects the reflected light, measuring its intensity modulated by blood volume changes. This varying intensity is converted into an electrical signal, creating the PPG waveform. Blood pressure signals are recorded using an invasive and continuous method, i.e., the catheter, and a non-invasive and discontinuous method, i.e., the blood pressure cuff. ECG is continuously recorded by placing electrodes on the patient’s chest. The Sainte-Justine University Hospital PICU utilizes a high-resolution research database (HRDB) [26, 27] that has been approved by the ethical committee. The HRDB links biomedical signals extracted from the different devices, displayed through patient monitors, to the electronic patient record continuously throughout their stay in the unit [28].
Between 2018 and 2022, 1571 patients met the inclusion criteria. For each patient, four physiological signals were extracted: ECG, PPG, blood pressure from the catheter, and blood pressure from the cuff. Each signal was extracted over 96 hours (4 days). Signal values are grouped together in a table with the date and time of acquisition. For the PPG signal, 640 values are acquired every 5 seconds, corresponding to a sampling frequency of 128 Hz. For blood pressure and ECG signals, 2560 values are acquired every 5 seconds, with a sampling frequency of 512 Hz. For the duration of the extraction, a fixed 30-second window of PPG signals will be used for further processing.
III-B Preprocessing
@@
    Filtering: each signal window is filtered using a band-pass Butterworth filter; the cut-off frequencies are 0.5 and 5 Hz, corresponding to a heart rate between 30 and 300 bpm. A forward-backward filtering is used to avoid phase distortions. The objective is to remove baseline wander and high-frequency noise.
@@
    Pulse segmentation: a function to find all local minima by comparing samples is used. The aim is to divide the preprocessed PPG signal into smaller segments or windows to detect the artifacts present for each pulse. In our case, a segment is a pulse. The size of each segment may vary depending on the characteristics of the PPG signal and the specific application of the signal pulses. A pulse is considered to lie between two minima.
@@
    Resampling: the duration of a cardiac cycle for children is between 0.3 and 1 second. A pulse represents a cardiac cycle. Therefore, not all pulses have the same number of samples. Each pulse is uniformly oversampled in time to contain 256 samples, corresponding to a heart cycle of 1s. A linear interpolation function [30] is used to create the missing points for each pulse. Linear interpolation is favored for signals due to its simplicity, computational efficiency, and ability to estimate values between known data points. It maintains signal continuity and linearity, making it suitable for signals with relatively smooth and linear variations.
@@
    Data transformation: each PPG pulse, essentially a waveform representing blood volume changes over time, can be represented as a data point in a column containing 256 values. These values are equally spaced points obtained using step 3 of the preprocessing. At the end of preprocessing, a vector of 256 points is obtained, representing a pulse of the PPG signal. The number of vectors depends on the number of pulses. This method allows us to work with PPG data in a structured manner suitable for various applications, from statistical analysis to machine learning.
Figure 2: Example of a 10s segment of a 30s raw PPG signal in the top image, filtered signal in the middle image, and segmented signal in the bottom image.
Fig. 2 shows the first 10 seconds of a raw PPG signal, when the signal has been filtered, and finally when the pulses have been segmented. The effect of the bandpass filter can be seen in the second figure. The filter has smoothed the signal by removing the extreme frequency components. The signal waveform is preserved, and the filter does not introduce resonances or significant ripples in the desired frequency range. Note that the first pulse has not been segmented. This is because the function could not detect the two minima that make up a pulse and, therefore, could not segment it. The signal does not start at the first low point of the pulse.
@@
III-C Dataset annotation
First, after preprocessing the data, the aim is to build a ground truth for future evaluation of the classification algorithms. To this end, one human expert annotated 10% of the database. Annotation is visual, comparing pulses with each other and binary classifying each pulse as good or as containing artifacts. To avoid involving another human expert, in consideration of time and specialist resources, we implemented an automated algorithm to handle additional annotations. This algorithm, acting as a surrogate expert, was developed to recheck the entirety of the 10% annotated data by the human expert, identifying similarities in the process. Employing a statistical approach, the algorithm determines if the values of a given pulse lie within standard parameters or deviate from the norm. We subsequently cross-validated the algorithm’s annotations against those from the professional expert to determine the algorithm’s accuracy. It was decided to annotate a maximum of 10% of the database and then use the LP algorithm to annotate the rest of the data.
III-C1 Expert labeling
PPG signals already segmented are presented to the expert. By analyzing each pulse, the expert classifies each pulse as artifact or artifact-free. A pulse is defined as artifact-free if its morphology is typical, i.e., if its characteristics - amplitude, width, shape - are the same as those of adjacent signals. A pulse is defined with artifacts if its characteristics differ from those of adjacent pulses (see fig. 3). To recheck the annotations of one expert, an algorithm that acts as a second expert was set up, allowing all impulses to be reannotated to see similarities. This algorithm uses a statistical approach to assess whether the statistical values of a pulse are normal or outside the norm.
III-C2 Statistical analysis
For each cardiac cycle, which corresponds to a pulse, if the waveform is similar, statistics such as skewness, standard deviation (std), and kurtosis are approximately constant for each cycle. It is, therefore, possible to detect motion artifacts by using the value of these statistics to differentiate a pulse without artifacts from a pulse with motion artifacts [16]. Skewness indicates the degree of asymmetry in the probability distribution of a random variable around its mean. It can take on positive, zero, negative, or undefined values, reflecting the shape and symmetry of the distribution. Kurtosis is the sharpened of the peak of a frequency-distribution curve, and standard deviation reflects the dispersion degree of a data set. If $X$ is considered a variable with $\mu$ and $\sigma$, the mean and standard deviation, respectively, statistical values are calculated as follows: